OpenStudy (anonymous):

Just Checking....The Q is: If it was 40°C outside, what was the temperature in degrees Fahrenheit? A. 85°F B. 98°F C. 104°F D. 132°F Is it C?

OpenStudy (anonymous):

c i think

OpenStudy (anonymous):

an easy way learned it was double the degrees C and subtract 10% then you alwasy add 32 to that

OpenStudy (anonymous):

the actual equation is F= 1.8(C) +32

OpenStudy (anonymous):

Thats what I Thought that's why I ask if it's C..40 Celsius is 104 Fahrenhiet Right?

OpenStudy (anonymous):

I Don't nessacarly need to do it that way. I believe I Got it right. And I Feel my ways easier to understand to me.

OpenStudy (anonymous):

1.8*40 =72 72+32 =104